what is the answer of 0 divided by zero
Answer by stanbon(75887)

what is the answer of 0 divided by zero
--------------------
Ans: undefined
=====================
Why?
6/3 = 2 because 3*2 = 6
-----------------------------
How about 0/0 = 9
True because 9*0 = 0
------------------------
How about 0/0 = 2
True because 2*0 = 0
--------
You see 0/0 could be any number.
Therefore 0/0 is "undefined".
==================================
Cheers,
Stan H.